# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

### v1

@@ 2.1. Notations
Regarding the topological degree $Q\in\mathbb{Z}$ defined in (1.2), through the paper we mainly consider the case $Q>0$ When $Q<0$, the same results follow by replacing the multi-soliton $m_{\xi_{1},\dots,\xi_{k}}=m_{\xi_{1}}+\dots+m_{\xi_{k}}$, consisting of kinks, with $m_{\xi_{1},\dots,\xi_{k}}^{-}=m_{\xi_{1}}^{-}+\dots+m_{\xi_{k}}^{-}$, consisting of anti-kinks.
@@ 2.2. Brownian bridges and homotopy-class representatives
$B_{\varepsilon,(-L_{\varepsilon},L_{\varepsilon})}^{0,2\pi Q}(x)=\frac{\pi Q}{L_{\varepsilon}}(x+L_{\varepsilon})+\sqrt{\varepsilon}B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x),$ (2.1)
@@
$\mathbb{E}\Big[\sqrt{\varepsilon}B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x_{1})\sqrt{\varepsilon}B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x_{2})\Big]=\frac{\varepsilon}{2L_{\varepsilon}}\big((x_{1}+L_{\varepsilon})(L_{\varepsilon}-x_{2})\wedge(x_{2}+L_{\varepsilon})(L_{\varepsilon}-x_{1})\big)$ (2.2)
@@
$e_{n}(x)=\frac{1}{\sqrt{L_{\varepsilon}}}\sin\Big(\frac{n\pi(x+L_{\varepsilon})}{2L_{\varepsilon}}\Big)$ (2.3)
@@
$B_{\varepsilon,(-L_{\varepsilon},L_{\varepsilon})}^{0,2\pi Q}(x)=\frac{\pi Q}{L_{\varepsilon}}(x+L_{\varepsilon})+\sqrt{\varepsilon}\sum_{n=1}^{\infty}\frac{2\sqrt{L_{\varepsilon}}g_{n}}{\pi n}\sin\Big(\frac{n\pi(x+L_{\varepsilon})}{2L_{\varepsilon}}\Big),$ (2.4)
@@ 2.3. Boué-Dupuis formula
$Y(t)=\sum_{n\geq 1}\frac{B_{n}(t)}{\sqrt{\lambda_{n}}}e_{n}(x)=\sum_{n=1}^{\infty}\frac{2\sqrt{L_{\varepsilon}}B_{n}(t)}{\pi n}\sin\Big(\frac{n\pi(x+L_{\varepsilon})}{2L_{\varepsilon}}\Big),$ (2.5)
@@
$\Theta(t):=\int_{0}^{t}(-\partial_{x}^{2})^{-\frac{1}{2}}\theta(s)ds.$ (2.6)
@@
$\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y(t)|^{2}dx\bigg]$ $=\frac{2}{3}t\cdot\varepsilon L_{\varepsilon}^{2},$ (2.7)
$\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y(t)|dx\bigg]$ $=\frac{\sqrt{\pi t}}{2}\cdot\varepsilon^{\frac{1}{2}}L_{\varepsilon}^{\frac{3}{2}}.$ (2.8)
@@
$\|\Theta(1)\|_{\dot{H}^{1}_{x}}^{2}\leq\int_{0}^{1}\|\dot{\Theta}(t)\|_{\dot{H}^{1}_{x}}^{2}dt=\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt,$ (2.9)
where $\dot{\Theta}(t)=\langle\nabla\rangle^{-1}\theta(t)$.
@@
Part (ii) follows from Minkowski’s and Cauchy-Schwarz’s inequalities. We first prove (2.7). Since $\{e_{n}\}_{n\geq 1}$ is an $L^{2}$-orthonormal basis of $L^{2}([-L_{\varepsilon},L_{\varepsilon}])$, we have $\|e_{n}\|_{L^{2}([-L_{\varepsilon},L_{\varepsilon}])}=1$ for all $n\geq 1$. This implies
$\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y(t)|^{2}dx\bigg]$ $=\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\bigg|\sqrt{\varepsilon}\sum_{n\geq 1}\frac{B_{n}(t)}{\sqrt{\lambda_{n}}}e_{n}(x)\bigg|^{2}dx\bigg]=t\varepsilon\sum_{n\geq 1}\frac{1}{\lambda_{n}}$
$=t\varepsilon(2L_{\varepsilon})^{2}\sum_{n\geq 1}\frac{1}{(n\pi)^{2}}=\frac{2}{3}t\cdot\varepsilon L_{\varepsilon}^{2},$
where we used $\lambda_{n}=(\tfrac{n\pi}{2L_{\varepsilon}})^{2}$. We now prove (2.8). From $Y(t)=\sum_{n\geq 1}\frac{B_{n}(t)}{\sqrt{\lambda_{n}}}e_{n}(x)$, (2.1), (2.4), and (2.2), we have
$\mathbb{E}|Y(t,x)|^{2}=t\sum_{n\geq 1}\frac{1}{\lambda_{n}}e_{n}(x)^{2}=t\mathbb{E}\Big[B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x)B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x)\Big]=t\frac{(x+L_{\varepsilon})(L_{\varepsilon}-x)}{2L_{\varepsilon}}.$
This implies that for every fixed $t$ and $x$, $\sqrt{\varepsilon}Y(t,x)\sim\mathcal{N}(0,\varepsilon t\tfrac{(x+L_{\varepsilon})(L_{\varepsilon}-x)}{2L_{\varepsilon}})$. Recall that for a centered Gaussian $Z\sim\mathcal{N}(0,\sigma^{2})$, $\mathbb{E}|Z|=\sigma\sqrt{\frac{2}{\pi}}$. Therefore,
$\mathbb{E}|\sqrt{\varepsilon}Y(t,x)|=\sqrt{\frac{2}{\pi}}\sqrt{\varepsilon t\frac{(x+L_{\varepsilon})(L_{\varepsilon}-x)}{2L_{\varepsilon}}}=\frac{\sqrt{\varepsilon t}}{\sqrt{\pi L_{\varepsilon}}}\sqrt{L_{\varepsilon}^{2}-x^{2}}.$ (2.10)
Using (2.10), we have
$\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y(t)|dx\bigg]=\frac{\sqrt{\varepsilon t}}{\sqrt{\pi L_{\varepsilon}}}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\sqrt{L_{\varepsilon}^{2}-x^{2}}dx=\frac{\sqrt{\varepsilon t}}{\sqrt{\pi L_{\varepsilon}}}\frac{\pi L_{\varepsilon}^{2}}{2}=\frac{\sqrt{\pi}}{2}\varepsilon^{\frac{1}{2}}t^{\frac{1}{2}}L_{\varepsilon}^{\frac{3}{2}}.$
@@ 3.2. Multi-topological solitons
$m_{\xi_{1},\dots,\xi_{k}}$ $=\sum_{j=1}^{k}m(\cdot-\xi_{j})=\sum_{j=1}^{k}m_{\xi_{j}}$
$m_{\xi_{1},\dots,\xi_{k}}^{-}$ $=\sum_{j=1}^{k}m^{-}(\cdot-\xi_{j})=\sum_{j=1}^{k}m^{-}_{\xi_{j}}.$
@@
Let $Q\in\mathbb{Z}$ with $|Q|=k$. There exists $0<\eta\ll 1$ sufficiently small such that
$\|\nabla E(m_{\xi_{1},\dots,\xi_{k}})\|_{L^{2}(\mathbb{R})}\leq C^{\nabla}\frac{e^{-(1-\eta)d}}{1-e^{-(1-\eta)d}}$
uniformly in $\xi_{1},\dots,\xi_{k}$ where $d:=\min_{i\neq j}|\xi_{i}-\xi_{j}|$ and $C^{\nabla}=(\frac{1}{2\eta}e^{\eta-1})^{1/2}Q$.
@@
$\|\nabla E(m_{\xi_{1},\dots,\xi_{k}})\|_{L^{2}(\mathbb{R})}\leq(\frac{1}{2\eta}e^{\eta-1})^{1/2}\sum_{i<j}e^{-(1-\eta)|\xi_{i}-\xi_{j}|}.$
One could estimate $\sum_{i<j}$ by $Q^{2}$, but to be slightly sharper, we assume without loss of generality that $\xi_{1}\leq\cdots\leq\xi_{k}$. Then there are Q1Q-1 terms of the form $e^{-(1-\eta)|\xi_{i}-\xi_{i+1}|}$, and Q2Q-2 terms of the form $e^{-(1-\eta)|\xi_{i}-\xi_{i+2}|}$ which can be written as $e^{-(1-\eta)|\xi_{i}-\xi_{i+1}|-(1-\eta)|\xi_{i+1}-\xi_{i+2}|}$, etc. Thus
$\|\nabla E(m_{\xi_{1},\dots,\xi_{k}})\|_{L^{2}(\mathbb{R})}\leq(\frac{1}{2\eta}e^{\eta-1})^{1/2}\sum_{k=1}^{Q-1}(Q-k)e^{-(1-\eta)kd}\leq(\frac{1}{2\eta}e^{\eta-1})^{1/2}Q\frac{e^{-(1-\eta)d}}{1-e^{-(1-\eta)d}}.$
@@
$\textup{dist}(\phi,\mathcal{M}^{Q}):=\inf_{\xi_{1},\dots,\xi_{Q}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}(\mathbb{R})}\geq\delta>0,$ (3.4)
@@
$\phi_{j}(x)=\begin{cases}\phi(x)\quad&x\in(t_{j-1},t_{j})\\ 2\pi(j-1)\quad&x<t_{j-1}\\ 2\pi j\quad&x>t_{j}.\end{cases}$ (3.5)
@@
$\geq|Q|E_{\textup{kink}}+c\sum_{j=1}^{Q}\inf_{\xi\in\mathbb{R}}\|(\phi_{j}-2\pi(j-1))-m_{\xi}\|_{L^{2}(\mathbb{R})}^{2}.$ (3.7)
@@
$\inf_{\xi_{1},\dots,\xi_{Q}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}(\mathbb{R})}^{2}\leq c\sum_{j=1}^{Q}d_{j}^{2},$ (3.8)
@@
$\textup{dist}(\phi,\mathcal{M}^{Q}):=\inf_{\xi_{1},\dots,\xi_{|Q|}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}^{-}\|_{L^{2}(\mathbb{R})}\geq\delta>0,$
where $m_{\xi_{1},\dots,\xi_{k}}^{-}=m^{-}(\cdot-\xi_{1})+\dots+m^{-}(\cdot-\xi_{|Q|})$.
@@
Let $Q\in\mathbb{Z}$ with $Q\geq 2$, and let $d>0$ be a large constant. Then there exists $c>0$ such that if $\phi\in\mathcal{C}_{Q}$ with $\phi(-\infty)=0$ and $\phi(\infty)=2\pi Q$ satisfies
$\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})=\inf_{\begin{subarray}{c}\xi_{1},\dots,\xi_{Q}\in\mathbb{R}\\ \min_{i\neq j}|\xi_{i}-\xi_{j}|<d\end{subarray}}\|\phi-m_{\xi_{1},\dots,\xi_{Q}}\|_{L^{2}(\mathbb{R})}<\delta$ (3.10)
where $\frac{1}{2}e^{-d}\geq c\cdot\delta$, then
@@
Since $\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta$, we can find $\xi_{1}^{*},\dots,\xi_{Q}^{*}$ such that
$\min_{i\neq j}|\xi_{i}^{*}-\xi_{j}^{*}|<d\quad\textup{and}\quad\|\phi-m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\|_{L^{2}(\mathbb{R})}\leq\delta.$ (3.11)
Therefore, from Lemma 3.12, we obtain
$E(\phi)$ $=\big(E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\big)+E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$
$\geq\big(E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\big)+|Q|E_{\textup{kink}}+c\cdot e^{-d}.$ (3.12)
By denoting $\eta(x):=\phi(x)-m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}(x)$, we write
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$ $=\int\partial_{x}m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\partial_{x}\eta dx+\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\eta|^{2}dx$
$\hphantom{X}+\int_{\mathbb{R}}(1-\cos(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}+\eta))dx-\int_{\mathbb{R}}(1-\cos m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})dx.$ (3.13)
Use Taylor on the cosine pointwise
$\cos(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}(x)+\eta(x))$
$=\cos m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}(x)-\sin m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}(x)\cdot\eta(x)-\frac{1}{2}\cos\big(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}(x)+\theta(x)\eta(x)\big)\cdot\eta(x)^{2}$ (3.14)
for some $\theta(x)\in(0,1)$. By plugging (3.14) into (3.13), we obtain
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$ $=\int\partial_{x}m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\partial_{x}\eta dx+\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\eta|^{2}dx$
$\hphantom{X}+\int_{\mathbb{R}}\sin m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\cdot\eta dx+\frac{1}{2}\int_{\mathbb{R}}\cos\big(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}+\theta\eta\big)\cdot\eta^{2}dx.$
@@
$E(\phi)$ $-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$
$=\int_{\mathbb{R}}\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\cdot\eta dx+\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\eta|^{2}dx+\frac{1}{2}\int_{\mathbb{R}}\cos\big(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}+\theta\eta\big)\cdot\eta^{2}dx,$
where $\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})=-\partial_{x}^{2}m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}+\sin m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}$. Using $\|\eta\|_{L^{2}(\mathbb{R})}=\|\phi-m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\|_{L^{2}(\mathbb{R})}\leq\delta$ and $\cos\geq-1$,
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$ $\geq-\|\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\|_{L^{2}(\mathbb{R})}\cdot\delta-\frac{1}{2}\|\eta\|_{L^{2}(\mathbb{R})}^{2}$
$\geq-\|\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\|_{L^{2}(\mathbb{R})}\cdot\delta-\frac{1}{2}\delta^{2}.$ (3.15)
From (3.3), we have $\|\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\|_{L^{2}(\mathbb{R})}\lesssim e^{-(1-\eta)\min_{i\neq j}|\xi_{i}^{*}-\xi_{j}^{*}|}$. This implies that
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\geq-c\delta-\frac{1}{2}\delta^{2}\geq-c_{1}\delta$ (3.16)
for all $0<\delta\leq\delta_{0}$, where $\delta_{0}>0$ is sufficiently small and $c_{1}>0$. Combining (3.16) and (3.12) yields
$E(\phi)\geq|Q|E_{\textup{kink}}+c\cdot e^{-d}-c_{1}\delta\geq|Q|E_{\textup{kink}}+\frac{c}{2}\cdot e^{-d}$ (3.17)
under the condition $\frac{c}{2}e^{-d}\geq c_{1}\delta$.
@@
Let $Q\in\mathbb{Z}$ with $Q\geq 2$. For $d>0$ sufficiently large, there exists $c>0$ such that for all $\xi_{1},\dots,\xi_{Q}\in\mathbb{R}$ with $\min_{i\neq j}|\xi_{i}-\xi_{j}|<d$, one has
@@
Now we need to prove that $\int_{\mathbb{R}}F_{Q}(\xi_{1},\cdots,\xi_{Q},x)dx\geq ce^{-\dot{d}}$. To this end we show the following representation of $F_{Q}$
@@
$=\frac{1}{2}\sum_{e}\Big(\cos((1-e_{1})m_{i_{1}}+\cdots+(1-e_{n})m_{i_{n}})-\cos((1+e_{1})m_{i_{1}}+\cdots+(1+e_{n})m_{i_{n}})\Big)\prod_{j=1}^{n}e_{j}$
$=-\sum_{e}\cos((1+e_{1})m_{i_{1}}+\cdots+(1+e_{n})m_{i_{n}})\Big)\prod_{j=1}^{n}e_{j}$ (3.20)
where the last step is by a change of variables $e_{j}\to-e_{j}$, and similarly for $n$ even,

### v2

@@ 2.1. Notations
Regarding the topological degree $Q\in\mathbb{Z}$ defined in (1.2), through the paper we mainly consider the case $Q>0$. When $Q<0$, the same results follow by replacing the multi-soliton $m_{\xi_{1},\dots,\xi_{k}}=m_{\xi_{1}}+\dots+m_{\xi_{k}}$, consisting of kinks, with $m_{\xi_{1},\dots,\xi_{k}}^{-}=m_{\xi_{1}}^{-}+\dots+m_{\xi_{k}}^{-}$, consisting of anti-kinks.
@@ 2.2. Brownian bridges and homotopy-class representatives
(2.1) $B_{\varepsilon,(-L_{\varepsilon},L_{\varepsilon})}^{0,2\pi Q}(x)=\frac{\pi Q}{L_{\varepsilon}}(x+L_{\varepsilon})+\sqrt{\varepsilon}B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x),$
@@
(2.2) $\mathbb{E}\Big[\sqrt{\varepsilon}B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x_{1})\sqrt{\varepsilon}B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x_{2})\Big]=\frac{\varepsilon}{2L_{\varepsilon}}\big((x_{1}+L_{\varepsilon})(L_{\varepsilon}-x_{2})\wedge(x_{2}+L_{\varepsilon})(L_{\varepsilon}-x_{1})\big)$
@@
(2.3) $e_{n}(x)=\frac{1}{\sqrt{L_{\varepsilon}}}\sin\Big(\frac{n\pi(x+L_{\varepsilon})}{2L_{\varepsilon}}\Big)$
@@
(2.4) $B_{\varepsilon,(-L_{\varepsilon},L_{\varepsilon})}^{0,2\pi Q}(x)=\frac{\pi Q}{L_{\varepsilon}}(x+L_{\varepsilon})+\sqrt{\varepsilon}\sum_{n=1}^{\infty}\frac{2\sqrt{L_{\varepsilon}}g_{n}}{\pi n}\sin\Big(\frac{n\pi(x+L_{\varepsilon})}{2L_{\varepsilon}}\Big),$
@@ 2.3. Boué-Dupuis formula
(2.5) $Y(t,x)=\sum_{n\geq 1}\frac{B_{n}(t)}{\sqrt{\lambda_{n}}}e_{n}(x)=\sum_{n=1}^{\infty}\frac{2\sqrt{L_{\varepsilon}}B_{n}(t)}{\pi n}\sin\Big(\frac{n\pi(x+L_{\varepsilon})}{2L_{\varepsilon}}\Big),$
@@
(2.6) $\Theta(t):=\int_{0}^{t}(-\partial_{x}^{2})^{-\frac{1}{2}}\theta(s)ds.$
@@
(2.7) $\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y(t)|^{2}dx\bigg]$ $=\frac{2}{3}t\cdot\varepsilon L_{\varepsilon}^{2},$
(2.8) $\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y(t)|dx\bigg]$ $=\frac{\sqrt{\pi t}}{2}\cdot\varepsilon^{\frac{1}{2}}L_{\varepsilon}^{\frac{3}{2}}.$
@@
(2.9) $\|\Theta(1)\|_{\dot{H}^{1}_{x}}^{2}\leq\int_{0}^{1}\|\dot{\Theta}(t)\|_{\dot{H}^{1}_{x}}^{2}dt=\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt,$
where $\dot{\Theta}(t)=(-\partial_{x}^{2})^{-\frac{1}{2}}\theta(t)$.
@@
For Part (i), (2.7) follows immediately from (2.5), Parseval’s identity, and $\sum_{n\geq 1}\frac{1}{n^{2}}=\frac{\pi^{2}}{6}$. To prove (2.8), by (2.5), (2.1), (2.4), and (2.2), we have
$\mathbb{E}|Y(t,x)|^{2}=t\sum_{n\geq 1}\frac{1}{\lambda_{n}}e_{n}(x)^{2}=t\mathbb{E}\Big[(B_{(-L_{\varepsilon},L_{\varepsilon})}^{0,0}(x))^{2}\Big]=t\frac{(x+L_{\varepsilon})(L_{\varepsilon}-x)}{2L_{\varepsilon}}.$
Namely $\sqrt{\varepsilon}Y(t,x)\sim\mathcal{N}(0,\varepsilon t\tfrac{(x+L_{\varepsilon})(L_{\varepsilon}-x)}{2L_{\varepsilon}})$. Recall that for a centered Gaussian $Z\sim\mathcal{N}(0,\sigma^{2})$, $\mathbb{E}|Z|=\sigma\sqrt{\frac{2}{\pi}}$. Therefore, (2.8) follows by computing $\frac{\sqrt{\varepsilon t}}{\sqrt{\pi L_{\varepsilon}}}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\sqrt{L_{\varepsilon}^{2}-x^{2}}\,dx$.
Part (ii) follows from Minkowski and Cauchy-Schwarz inequalities.
@@ 3.2. Multi-topological solitons
$m_{\xi_{1},\dots,\xi_{k}}=\sum_{j=1}^{k}m(\cdot-\xi_{j})=\sum_{j=1}^{k}m_{\xi_{j}},\qquad m_{\xi_{1},\dots,\xi_{k}}^{-}=\sum_{j=1}^{k}m^{-}(\cdot-\xi_{j})=\sum_{j=1}^{k}m^{-}_{\xi_{j}}.$
@@
Let $Q\in\mathbb{Z}$ with $|Q|=k$. For $0<\eta\ll 1$ sufficiently small,
$\|\nabla E(m_{\xi_{1},\dots,\xi_{k}})\|_{L^{2}(\mathbb{R})}\leq C_{\eta}e^{-(1-\eta)d},$
uniformly in $\xi_{1},\dots,\xi_{k}$ where $d:=\min_{i\neq j}|\xi_{i}-\xi_{j}|$ and $C_{\eta}=O(\eta^{-1/2})$.
@@
$\|\nabla E(m_{\xi_{1},\dots,\xi_{k}})\|_{L^{2}(\mathbb{R})}\leq\Big(\frac{1}{2\eta}e^{\eta-1}\Big)^{1/2}\sum_{i<j}e^{-(1-\eta)|\xi_{i}-\xi_{j}|}.$
Thus the bound follows. ∎
@@
(3.4) $\textup{dist}(\phi,\mathcal{M}^{Q}):=\inf_{\xi_{1},\dots,\xi_{Q}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}(\mathbb{R})}\geq\delta>0,$
@@
(3.5) $\phi_{j}(x)=\begin{cases}\phi(x)\quad&x\in(t_{j-1},t_{j})\\ 2\pi(j-1)\quad&x<t_{j-1}\\ 2\pi j\quad&x>t_{j}.\end{cases}$
@@
(3.7) $\geq|Q|E_{\textup{kink}}+c\sum_{j=1}^{Q}\inf_{\xi\in\mathbb{R}}\|(\phi_{j}-2\pi(j-1))-m_{\xi}\|_{L^{2}(\mathbb{R})}^{2}.$
@@
(3.8) $\inf_{\xi_{1},\dots,\xi_{Q}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}(\mathbb{R})}^{2}\leq c\sum_{j=1}^{Q}d_{j}^{2},$
@@
$\textup{dist}(\phi,\mathcal{M}^{Q}):=\inf_{\xi_{1},\dots,\xi_{|Q|}\in\mathbb{R}}\|\phi-m_{\xi_{1},\dots,\xi_{k}}^{-}\|_{L^{2}(\mathbb{R})}\geq\delta>0.$
3.3. Collision regime
@@
Let $Q\in\mathbb{Z}$ with $Q\geq 2$, and let $d>0$ be a large constant. Then there exist $c>0$ and $C>0$ such that if $\phi\in\mathcal{C}_{Q}$ with $\phi(-\infty)=0$ and $\phi(\infty)=2\pi Q$ satisfies
(3.10) $\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})=\inf_{\begin{subarray}{c}\xi_{1},\dots,\xi_{Q}\in\mathbb{R}\\ \min_{i\neq j}|\xi_{i}-\xi_{j}|<d\end{subarray}}\|\phi-m_{\xi_{1},\dots,\xi_{Q}}\|_{L^{2}(\mathbb{R})}<\delta$
where $\delta\leq c\,e^{-d/2}$, then
@@
Since $\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta$, we can find $m^{*}:=m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}$ such that
(3.11) $s:=\min_{i\neq j}|\xi_{i}^{*}-\xi_{j}^{*}|<d\quad\textup{and}\quad\|\phi-m^{*}\|_{L^{2}(\mathbb{R})}\leq\delta.$
Denoting $\eta(x):=\phi(x)-m^{*}(x)$, we write
$E(\phi)-E(m^{*})$ $=\int\partial_{x}m^{*}\partial_{x}\eta dx+\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\eta|^{2}dx$
$\hphantom{X}+\int_{\mathbb{R}}(1-\cos(m^{*}+\eta))dx-\int_{\mathbb{R}}(1-\cos m^{*})dx$
$=\int\partial_{x}m^{*}\partial_{x}\eta dx+\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\eta|^{2}dx$
$\hphantom{X}+\int_{\mathbb{R}}\sin m^{*}\cdot\eta+\frac{1}{2}\cos\big(m^{*}+\theta\eta\big)\cdot\eta^{2}\,dx$
where we Taylor expanded $\cos(m^{*}(x)+\eta(x))$ in $\eta(x)$ and $\theta(x)\in(0,1)$.
@@
$E(\phi)$ $-E(m^{*})=\int_{\mathbb{R}}\nabla E(m^{*})\cdot\eta dx+\frac{1}{2}\int_{\mathbb{R}}|\partial_{x}\eta|^{2}dx+\frac{1}{2}\int_{\mathbb{R}}\cos\big(m^{*}+\theta\eta\big)\cdot\eta^{2}dx,$
where $\nabla E(m^{*})=-\partial_{x}^{2}m^{*}+\sin m^{*}$. Using $\|\eta\|_{L^{2}(\mathbb{R})}=\|\phi-m^{*}\|_{L^{2}(\mathbb{R})}\leq\delta$ and $\cos\geq-1$,
(3.12) $E(\phi)-E(m^{*})\geq-\|\nabla E(m^{*})\|_{L^{2}(\mathbb{R})}\cdot\delta-\frac{1}{2}\delta^{2}.$
By Lemma 3.12 below,
(3.13) $E(m^{*})\geq|Q|E_{\textup{kink}}+C_{1}e^{-s},$
where $s=\min_{i\neq j}|\xi_{i}^{*}-\xi_{j}^{*}|$ is defined in (3.11). By Lemma 3.6,
(3.14) $\|\nabla E(m^{*})\|_{L^{2}(\mathbb{R})}\leq C_{\eta}e^{-(1-\eta)s}.$
Recall that $d$ is fixed, while $s$ depends on $\phi$. In the following we consider two cases.
Case 1: “Strong Collision” regime $s\leq d/2$. In this regime, the solitons are very close, leading to a large energy surplus. (3.13) implies
$E(m^{*})\geq|Q|E_{\textup{kink}}+C_{1}e^{-d/2}.$
For the gradient term we simply bound $\|\nabla E(m^{*})\|_{L^{2}(\mathbb{R})}\leq C_{3}$. Substituting these into (3.12), we have
$E(\phi)-|Q|E_{\textup{kink}}\geq C_{1}e^{-d/2}-C_{3}\delta-\frac{1}{2}\delta^{2}.$
Since $\delta\leq c\,e^{-d/2}$, for $c>0$ sufficiently small, the term $C_{1}e^{-d/2}$ dominates the other two terms on the right-hand side, so we have
$E(\phi)-|Q|E_{\textup{kink}}\geq\frac{C_{1}}{2}e^{-d/2}\geq Ce^{-d}.$
Case 2: “Weak collision” regime $\frac{d}{2}<s<d$. In this regime, the separation is larger, so the energy gap $e^{-s}$ is smaller, but the gradient is also exponentially small. From (3.12), (3.13), and (3.14) we have
(3.15) $E(\phi)-|Q|E_{\textup{kink}}\geq C_{1}e^{-s}-C_{\eta}e^{-(1-\eta)s}\delta-\frac{1}{2}\delta^{2}.$
Note that for $c>0$ small enough,
$\frac{1}{2}\delta^{2}\leq\frac{1}{2}c^{2}e^{-d}\leq\frac{C_{1}}{8}e^{-s},$
where we used $\delta\leq ce^{-d/2}$, and by choosing for instance $\eta=1/4$,
$C_{\eta}e^{-(1-\eta)s}\delta\leq C_{\eta}e^{-s+\eta d}\cdot c\,e^{-d/2}\leq\frac{C_{1}}{8}e^{-s}$
so the first term $C_{1}e^{-s}$ in (3.15) again dominates the other two terms on the right-hand side, and
$E(\phi)-|Q|E_{\textup{kink}}\geq Ce^{-s}\geq Ce^{-d}$
for some $C>0$.
@@
Let $Q\in\mathbb{Z}$ with $Q\geq 2$. For $d>0$ sufficiently large, there exists $c>0$ such that for all $\xi_{1},\dots,\xi_{Q}\in\mathbb{R}$ with $\min_{i\neq j}|\xi_{i}-\xi_{j}|=d$, one has
@@
Now we need to prove that $\int_{\mathbb{R}}F_{Q}(\xi_{1},\cdots,\xi_{Q},x)dx\geq c\cdot e^{-d}$. To this end we show the following representation of $F_{Q}$
@@
$=\frac{1}{2}\sum_{e}\Big(\cos((1-e_{1})m_{i_{1}}+\cdots+(1-e_{n})m_{i_{n}})$
$\qquad\qquad\qquad-\cos((1+e_{1})m_{i_{1}}+\cdots+(1+e_{n})m_{i_{n}})\Big)\prod_{j=1}^{n}e_{j}$
(3.18) $=-\sum_{e}\cos((1+e_{1})m_{i_{1}}+\cdots+(1+e_{n})m_{i_{n}})\Big)\prod_{j=1}^{n}e_{j}$
where the last step is by a change of variables $e_{j}\to-e_{j}$. Similarly for $n$ even,